# StructCore: Structure-Aware Image-Level Scoring for Training-Free Anomaly Detection thanks: Citation: StructCore: Structure-Aware Image-Level Scoring for Training-Free Anomaly Detection

### v1

@@ Abstract
Max pooling is the de facto standard for converting anomaly score maps into image-level decisions in memory-bank-based unsupervised anomaly detection (UAD). However, because it relies on a single extreme response, it discards most information about how anomaly evidence is distributed and structured across the image, often causing normal and anomalous scores to overlap. We propose StructCore, a training-free, structure-aware image-level scoring method that goes beyond max pooling. Given an anomaly score map, StructCore computes a low-dimensional structural descriptor $\phi(S)$ that captures distributional and spatial characteristics, and refines image-level scoring via a Mahalanobis-based calibration fitted on train-good samples, without modifying pixel-level localization. StructCore achieves image-level AUROC scores of 99.6% on MVTec AD and 98.4% on VisA, demonstrating robust image-level anomaly detection by exploiting structural signatures missed by max pooling.
@@ 1 Introduction
Unsupervised Anomaly Detection (UAD) is a core component of industrial visual inspection, where anomalies are rare, diverse, and costly to annotate. In many production settings, only normal (train-good) samples are available for training, and the primary operational requirement is a reliable image-level accept/reject decision under strict latency and memory budgets [4]. While pixel-level localization is important, the final image-level decision rule often determines practical system reliability.
@@
This perspective is further motivated by a counter-intuitive observation in recent ViT-based reconstruction-style methods such as Dinomaly [17]. By leveraging global context in Vision Transformers, these methods can produce highly refined normal feature maps and mitigate the classical “identity mapping” issue, which might suggest that image-level detection should approach a performance ceiling. Yet, in practice, simple max pooling can still fail to provide the expected separation in certain cases. We interpret this as a structural gap: even when anomaly maps are high-quality, decisive evidence may reside not in a single extreme response but in the collective distributional and spatial structure of the map, which max pooling inherently ignores.
Motivated by this, we propose StructCore, a training-free, structure-aware image-level scoring module that goes beyond max pooling. Given an anomaly score map $S$, StructCore computes a compact, low-dimensional structural descriptor $\phi(S)$ that summarizes distributional and spatial signatures of the map. Using only train-good samples, we estimate normal statistics of $\phi(S)$ and compute a Mahalanobis-style (diagonal standardization-based) calibration score to refine image-level decisions [22]. Importantly, StructCore refines image-level scoring without modifying pixel-level localization: patch scores and anomaly maps remain unchanged.
@@ 2 Related Work
Industrial visual anomaly detection (VAD) is commonly studied under a one-class protocol, where only normal (train-good) images are available during training and the goal is to detect and localize defects at test time. Benchmarks such as MVTec AD [4] and VisA [47] standardize evaluation by reporting both image-level detection and pixel-level localization. Despite the diversity of model families, a large portion of practical industrial pipelines share a consistent interface: they first produce an anomaly score map and then reduce it to a single scalar image score for accept/reject decisions. In this work, we target this interface explicitly. Instead of modifying pixel-level anomaly maps (and thus localization behavior), we focus on improving the final map-to-score decision rule at image level.
@@
Several works in anomaly detection implicitly revisit the map-to-score issue through alternative aggregation or calibration strategies. Unified frameworks may employ top-$k$ or top-ratio matching to construct image-level anomaly scores [2], and recent test-time re-scoring methods aim to reduce false positives caused by noisy peaks by recalibrating image-level decisions without retraining [24, 20]. Compared to these approaches, StructCore formalizes a simple and fully training-free principle: summarize the anomaly map using a low-dimensional structural descriptor and calibrate it using train-good statistics via a Mahalanobis-style distance [22]. This directly compensates for the information loss of max pooling while keeping the anomaly map unchanged.
@@ 3.4 Routed Memory Inference for MUAD
In a multi-category setting, exhaustive scoring against all category memory banks scales linearly with $|\mathcal{C}|$ and increases deployment cost. To avoid this, we use distance-based routing to bind each input to a single category and then query only the corresponding memory bank $\mathbf{M}_{\hat{c}}$.
For each category $c$, we construct a routing bank $\mathbf{R}_{c}\in\mathbb{R}^{K_{r}\times D}$ consisting of $K_{r}$ prototypes (default $K_{r}{=}64$). In a normalized space, we compute the nearest-prototype distance for each patch and aggregate over patches to obtain a routing score:
$d_{p,c}=\min_{r\in\mathbf{R}_{c}}\lVert\tilde{z}_{p}-\tilde{r}\rVert_{2}^{2},\qquad\mathrm{route}(c)=\frac{1}{P}\sum_{p=1}^{P}d_{p,c},$ (3)
@@ 3.6 StructCore: Structure-aware Image-level Scoring
We fit normal statistics of $\phi(S)$ using train-good samples only. Let $\mu\in\mathbb{R}^{3}$ and $\sigma\in\mathbb{R}^{3}$ denote the per-dimension mean and standard deviation of $\phi(S)$ over train-good anomaly maps. At test time, we compute a diagonal Mahalanobis (standardized Euclidean) distance:
@@ 4.1 Experimental Settings
We evaluate StructCore on two standard industrial UAD benchmarks, MVTec AD (15 categories) and VisA (12 categories). Following the one-class protocol, all components are fitted using only normal (train-good) samples, and evaluation is performed on test sets containing both normal and anomalous images.
@@
Feature extraction: We use a frozen DINOv2 ViT-B/14 backbone (patch size 14) [32]. Preprocessing: We follow the preprocessing protocol of Dinomaly [17]: images are resized to $448\times 448$ and center-cropped to $392\times 392$, producing a $28\times 28$ token grid. Base detector: We use a PatchCore-style memory-based detector with skip-layer fusion, coreset selection, and kNN retrieval. StructCore: StructCore operates on anomaly maps without gradient-based training and refines only the image-level score (the anomaly map is unchanged). We use the 3D descriptor $\phi(S)=\{\sigma_{S},\mathrm{top1\%},\mathrm{TV}\}$ (Sec. 3.6) and compute $D_{\mathrm{struct}}$ via a diagonal Mahalanobis (standardized Euclidean) distance. To avoid tuning, we use an automatic scale calibration $\lambda_{\mathrm{auto}}$ estimated from train-good statistics (Sec. 3.6).
@@
StructCore 3D $\phi(S)=\{\sigma_{S},\mathrm{top1\%},\mathrm{TV}\}$; diag Mahalanobis; auto-$\lambda$ (train-good statistics)
@@ 4.2 Comparison
Table 3: Performance (%) on MVTec-AD and VisA. Our StructCore results are reported with 1%) coreset. Best results are in bold.
@@ 4.3 Ablation Study
StructCore uses the hybrid weight $\lambda$ to control the contribution of the structural calibration term. We perform a sensitivity analysis over $\lambda\in\{0.05,0.10,0.15,0.20\}$ across all categories. While the optimal $\lambda$ may vary by category, the overall mean image-level AUROC remains essentially unchanged (around 99.6%) across these values, indicating that StructCore is not overly sensitive to $\lambda$ in this range. Accordingly, we use a single fixed default $\lambda{=}0.05$ in all experiments without category-wise tuning.
@@
Table 5 performs a controlled ablation by fixing the base detector and varying only $\phi(S)$. Each component alone yields a clear improvement over the base image score (e.g., $\{\sigma_{S}\}$: +0.89, $\{\mathrm{top1\%}\}$: +0.88), indicating that both global dispersion and tail evidence provide useful calibration signals. Combining components is consistently beneficial, and the full 3D descriptor $\{\sigma_{S},\mathrm{top1\%},\mathrm{TV}\}$ achieves the best performance (+0.99), supporting our design choice of a minimal yet effective structural basis.
Table 5: Ablation of $\phi(S)$ components on MVTec AD (mean I-AUROC, %). Base uses max pooling; Hybrid uses $S_{\mathrm{hyb}}=S_{\mathrm{base}}+D_{\mathrm{struct}}$ with. $\Delta$ denotes improvement over Base
@@
$\{\sigma_{S},\mathrm{top1\%}\}$ 98.68 99.66 +0.97
@@
$\{\mathrm{top1\%},\mathrm{TV}\}$ 98.68 99.59 +0.90
$\{\sigma_{S},\mathrm{top1\%},\mathrm{TV}\}$ 98.68 99.67 +0.99
@@
StructCore computes a structural score by measuring how atypical $\phi(S)$ is relative to train-good statistics. While we use a Mahalanobis-style distance by default, other distance functions are possible. To assess sensitivity to this choice, we compare several distances (e.g., $\ell_{1}$, $\ell_{2}$, cosine, and Chebyshev) as well as their standardized variants that normalize each dimension by the train-good standard deviation. Table 6 reports mean image-level AUROC (mean $\pm$ std over categories). Across both datasets, standardized distances perform consistently well, and Mahalanobis remains competitive, indicating that StructCore is not overly sensitive to the particular distance choice. Unless otherwise stated, we use the Mahalanobis-style distance for simplicity and for consistency with prior calibration literature.
@@ A.1 Detailed Analysis of Structural Components
The proposed StructCore significantly enhances detection in categories where the Base (max pooling) method typically struggles due to the lack of structural context. For instance, in Cable ($92.71\%\rightarrow 99.10\%$), Capsule ($95.81\%\rightarrow 98.40\%$), and Screw ($96.82\%\rightarrow 98.38\%$), the inclusion of structural features effectively captures anomalies that are missed by a single peak response. This confirms that structural statistics are particularly vital for defects that are subtle or spatially distributed. While individual components like global dispersion ($\{\sigma_{S}\}$) or tail-based statistics ($\{\mathrm{topk\_mean}\}$) alone provide substantial boosts over the baseline, the full combination ($\{\sigma_{S},\mathrm{topk\_mean},\mathrm{TV}\}$) achieves the highest mean AUROC of $99.67\%$. This indicates that global dispersion, tail concentration, and spatial roughness capture complementary aspects of anomaly maps, leading to a more robust decision boundary. For categories where the Base model already achieves near-perfect performance (e.g., Bottle, Hazelnut, Leather), the Hybrid scoring maintains $100.00\%$ accuracy. This demonstrates that the structural calibration is safe and does not induce negative transfer or degradation in simpler cases where max pooling is already sufficient.
@@
Category Base $\{\sigma_{S}\}$ $\{\mathrm{topk\_mean}\}$ $\{\mathrm{TV}\}$ $\{\sigma_{S},\mathrm{topk\_mean}\}$ $\{\sigma_{S},\mathrm{TV}\}$ $\{\mathrm{topk\_mean},\mathrm{TV}\}$ $\{\sigma_{S},\mathrm{topk\_mean},\mathrm{TV}\}$

### v2

# StructCore: Structure-Aware Image-Level Scoring for Anomaly Detection thanks: Citation: StructCore: Structure-Aware Image-Level Scoring for Anomaly Detection

@@ Abstract
Max pooling is the de facto standard for converting anomaly score maps into image-level decisions in memory-bank-based unsupervised anomaly detection (UAD). However, because it relies on a single extreme response, it discards most information about how anomaly evidence is distributed and structured across the image, often causing normal and anomalous scores to overlap. We propose StructCore, a training-free, structure-aware image-level scoring method that goes beyond max pooling. Given an anomaly score map, StructCore computes a low-dimensional structural descriptor $\phi(S)$ that captures distributional and spatial characteristics, and refines image-level scoring via a diagonal Mahalanobis calibration estimated from train-good samples, without modifying pixel-level localization. StructCore achieves image-level AUROC scores of 99.6% on MVTec AD and 98.4% on VisA, demonstrating robust image-level anomaly detection by exploiting structural signatures missed by max pooling.
@@ 1 Introduction
Unsupervised Anomaly Detection (UAD) is a core component of industrial visual inspection, where anomalies are rare, diverse, and costly to annotate. In many production settings, only normal samples are available for training, and the primary operational requirement is a reliable image-level accept/reject decision under strict latency and memory budgets [3]. While pixel-level localization is important, the final image-level decision rule often determines practical system reliability.
@@
This perspective is further motivated by a counterintuitive observation in recent ViT-based reconstruction-style methods such as Dinomaly [16]. By leveraging global context in Vision Transformers, these methods can produce highly refined normal feature maps and mitigate the classical “identity mapping” issue, which might suggest that image-level detection should approach a performance ceiling. Yet, in practice, simple max pooling can still fail to provide the expected separation in certain cases. We interpret this as a structural gap: even when anomaly maps are high-quality, decisive evidence may reside not in a single extreme response but in the collective distributional and spatial structure of the map, which max pooling inherently ignores.
Motivated by this, we propose StructCore, a training-free, structure-aware image-level scoring module that goes beyond max pooling. Given an anomaly score map $S$, StructCore computes a compact, low-dimensional structural descriptor $\phi(S)$ that summarizes distributional and spatial signatures of the map. Using only train-good samples, we estimate normal statistics of $\phi(S)$ and compute a diagonal Mahalanobis calibration score to refine image-level decisions [21]. Importantly, StructCore refines image-level scoring without modifying pixel-level localization: patch scores and anomaly maps remain unchanged.
@@ 2 Related Work
Industrial visual anomaly detection (VAD) is commonly studied under a one-class protocol, where only normal images are available during training and the goal is to detect and localize defects at test time. Benchmarks such as MVTec AD [3] and VisA [47] standardize evaluation by reporting both image-level detection and pixel-level localization. Despite the diversity of model families, a large portion of practical industrial pipelines share a consistent interface: they first produce an anomaly score map and then reduce it to a single scalar image score for accept/reject decisions. In this work, we target this interface explicitly. Instead of modifying pixel-level anomaly maps (and thus localization behavior), we focus on improving the final map-to-score decision rule at image level.
@@
Several works in anomaly detection implicitly revisit the map-to-score issue through alternative aggregation or calibration strategies. Unified frameworks may employ top-$k$ or top-ratio matching to construct image-level anomaly scores [23], and recent test-time re-scoring methods aim to reduce false positives caused by noisy peaks by recalibrating image-level decisions without retraining [24, 19]. Compared to these approaches, StructCore formalizes a simple and fully training-free principle: summarize the anomaly map using a low-dimensional structural descriptor and calibrate it using train-good statistics via a diagonal Mahalanobis distance [21]. This directly compensates for the information loss of max pooling while keeping the anomaly map unchanged.
@@ 3.4 Routed Memory Inference for MUAD
We consider multi-category UAD (MUAD), where a single deployed system must handle multiple categories $\mathcal{C}$ under a unified inference interface. In this setting, exhaustive scoring against all category memory banks scales linearly with $|\mathcal{C}|$ and increases deployment cost. To reduce this overhead, we use distance-based routing to bind each input to a single category and then query only the corresponding memory bank $\mathbf{M}_{\hat{c}}$. At test time, the category identity is not assumed to be given, so the system must both select a relevant category memory and compute the anomaly score.
For each category $c$, we construct a routing bank $\mathbf{R}_{c}\in\mathbb{R}^{K_{r}\times D}$ consisting of $K_{r}$ prototypes (default $K_{r}{=}64$). We $\ell_{2}$-normalize embeddings for routing; $\tilde{z}$ and $\tilde{r}$ denote normalized patch embeddings and prototypes, respectively. In this normalized space, we compute the nearest-prototype distance for each patch and aggregate over patches to obtain a routing score:
$d_{p,c}=\min_{r\in\mathbf{R}_{c}}\lVert\tilde{z}_{p}-\tilde{r}\rVert_{2}^{2},\qquad\mathrm{route}(c)=\frac{1}{P}\sum_{p=1}^{P}d_{p,c}.$ (3)
@@ 3.6 StructCore: Structure-aware Image-level Scoring
We fit normal statistics of $\phi(S)$ using train-good samples only. Let $\mu\in\mathbb{R}^{3}$ and $\sigma\in\mathbb{R}^{3}$ denote the per-dimension mean and standard deviation of $\phi(S)$ over train-good anomaly maps. At test time, we compute a diagonal Mahalanobis (i.e., standardized Euclidean) distance:
@@ 4.1 Experimental Settings
We evaluate StructCore on two standard industrial UAD benchmarks, MVTec AD (15 categories) and VisA (12 categories). Following the one-class protocol, all components are fitted using only train-good samples, and evaluation is performed on test sets containing both normal and anomalous images.
@@
Feature extraction: We use a frozen DINOv2 ViT-B/14 backbone (patch size 14) [32]. Preprocessing: We follow the preprocessing protocol of Dinomaly [16]: images are resized to $448\times 448$ and center-cropped to $392\times 392$, producing a $28\times 28$ token grid. Base detector: We use a PatchCore-style memory-based detector with skip-layer fusion, coreset selection, and kNN retrieval. StructCore: StructCore operates on anomaly maps without gradient-based training and refines only the image-level score (the anomaly map is unchanged). We use the 3D descriptor $\phi(S)=\{\sigma_{S},\mathrm{topk\_mean}_{r},\mathrm{TV}\}$ (Sec. 3.6) and compute $D_{\mathrm{struct}}$ via a diagonal Mahalanobis distance. To avoid tuning, we use an automatic scale calibration $\lambda_{\mathrm{auto}}$ estimated from train-good statistics (Sec. 3.6).
@@
StructCore 3D $\phi(S)=\{\sigma_{S},\mathrm{topk\_mean}_{r},\mathrm{TV}\}$; diagonal Mahalanobis; auto-$\lambda$ (train-good statistics)
@@ 4.2 Comparison
Table 3: Performance (%) on MVTec-AD and VisA. Our StructCore results are reported with a 1% coreset. Best results are in bold.
@@ 4.3 Ablation Study
StructCore uses the hybrid weight $\lambda$ to control the contribution of the structural calibration term. We perform a sensitivity analysis over a fixed $\lambda\in\{0.05,0.10,0.15,0.20\}$ across all categories. While the optimal $\lambda$ may vary by category, the overall mean image-level AUROC remains essentially unchanged (around 99.6%) across these values, indicating that StructCore is not overly sensitive to $\lambda$ in this range. Unless stated otherwise, we use the automatic weight $\lambda_{\mathrm{auto}}$ in all experiments to avoid any tuning.
@@
Table 5 performs a controlled ablation by fixing the base detector and varying only $\phi(S)$. Each component alone yields a clear improvement over the base image score (e.g., $\{\sigma_{S}\}$: +0.89, $\{\mathrm{topk\_mean}_{r}\}$: +0.88), indicating that both global dispersion and tail evidence provide useful calibration signals. Combining components is consistently beneficial, and the full 3D descriptor $\{\sigma_{S},\mathrm{topk\_mean}_{r},\mathrm{TV}\}$ achieves the best performance (+0.99), supporting our design choice of a minimal yet effective structural basis.
Table 5: Ablation of $\phi(S)$ components on MVTec AD (mean I-AUROC, %). Base uses max pooling; Hybrid uses $S_{\mathrm{hyb}}=S_{\mathrm{base}}+\lambda_{\mathrm{auto}}D_{\mathrm{struct}}$. $\Delta$ denotes the improvement over Base.
@@
$\{\sigma_{S},\mathrm{topk\_mean}_{r}\}$ 98.68 99.66 +0.97
@@
$\{\mathrm{topk\_mean}_{r},\mathrm{TV}\}$ 98.68 99.59 +0.90
$\{\sigma_{S},\mathrm{topk\_mean}_{r},\mathrm{TV}\}$ 98.68 99.67 +0.99
@@
StructCore computes a structural score by measuring how atypical $\phi(S)$ is relative to train-good statistics. While we use the diagonal Mahalanobis distance by default, other distance functions are possible. To assess sensitivity to this choice, we compare several distances (e.g., $\ell_{1}$, $\ell_{2}$, cosine, and Chebyshev) as well as their standardized variants that normalize each dimension by the train-good standard deviation. Table 6 reports mean image-level AUROC (mean $\pm$ std over categories). Across both datasets, standardized distances perform consistently well, and diagonal Mahalanobis remains competitive, indicating that StructCore is not overly sensitive to the particular distance choice. Unless otherwise stated, we use the diagonal Mahalanobis distance for simplicity and for consistency with prior calibration literature.
@@ A.1 Detailed Analysis of Structural Components
The proposed StructCore significantly enhances detection in categories where the Base (max pooling) method typically struggles due to the lack of structural context. For instance, in Cable ($92.71\%\rightarrow 99.10\%$), Capsule ($95.81\%\rightarrow 98.40\%$), and Screw ($96.82\%\rightarrow 98.38\%$), the inclusion of structural features effectively captures anomalies that are missed by a single peak response. This confirms that structural statistics are particularly vital for defects that are subtle or spatially distributed. While individual components like global dispersion ($\{\sigma_{S}\}$) or tail-based statistics ($\{\mathrm{topk\_mean}_{r}\}$) alone provide substantial boosts over the baseline, the full combination ($\{\sigma_{S},\mathrm{topk\_mean}_{r},\mathrm{TV}\}$) achieves the highest mean AUROC of $99.67\%$.
This indicates that global dispersion, tail concentration, and spatial roughness capture complementary aspects of anomaly maps, leading to a more robust decision boundary. For categories where the Base model already achieves near-perfect performance (e.g., Bottle, Hazelnut, Leather), the Hybrid scoring maintains $100.00\%$ accuracy. This demonstrates that the structural calibration is safe and does not induce negative transfer or degradation in simpler cases where max pooling is already sufficient.
@@
Category Base $\{\sigma_{S}\}$ $\{\mathrm{topk\_mean}_{r}\}$ $\{\mathrm{TV}\}$ $\{\sigma_{S},\mathrm{topk\_mean}_{r}\}$ $\{\sigma_{S},\mathrm{TV}\}$ $\{\mathrm{topk\_mean}_{r},\mathrm{TV}\}$ $\{\sigma_{S},\mathrm{topk\_mean}_{r},\mathrm{TV}\}$